OpenStudy (anonymous):

Which fraction is not equivalent to 8 -- 10 a 4 -- 5 b 16 -- 20 c 8 - -- 10 d -8 -- -10

OpenStudy (anonymous):

well is 8/10 equivalent to -8/10? if you want to check by cross-multiplying, 8*5=40 and 10*4=40, 40=40 so 8/10=4/5. But -8*10=-80, 10*8=80, and -80 does not equal 80.

OpenStudy (anonymous):

@mattvlaw to the question you asked i dont believe so

OpenStudy (anonymous):

yep, so that's your answer